OpenStudy (qqstory):

1/1-x + 1/1+radical x = 1/ 1-radical x

OpenStudy (qqstory):

\[1 / 1 - x + 1/ 1+ \sqrt{x} = 1 / 1 - \sqrt{x}\]

OpenStudy (alexwee123):

 is it that?